OpenStudy (anonymous):

What is the product of (n - 8)(n + 2)? am I right n^2+10n-16

OpenStudy (aravindg):

try again ?

OpenStudy (aravindg):

(x+a)(x+b)=x^2+(a+b)x+ab

OpenStudy (anonymous):

n^2-10n-16

OpenStudy (aravindg):

a+b =?

OpenStudy (anonymous):

will u do my problem u r confusing the hell out of me

OpenStudy (aravindg):

just add constant terms to get coefficient of x !

OpenStudy (aravindg):

-10 is not coeffecient of x

OpenStudy (anonymous):

n^2 - 8n + 2n - 16 = n^2 -6n -16